OpenStudy (anonymous):

Need help really fasst!!!!! How many units apart are the points (-8, -8) and (4, 8)?

OpenStudy (thecatman):

got a pic?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[\sqrt{(4 + 8)^{2} + (8 + 8)^{2}}\] \[\sqrt{12^{2} + 16^{2}}\] \[\sqrt{144 + 256}\] \[\sqrt{400}\] \[20\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

r u sure about this answer?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

ok